# Dynamical Constraints on the Vertical Structure of Jupiter’s Polar Cyclones

### v1

@@ Abstract
Jupiter’s poles feature striking polygons of cyclones that drift westward over time—a motion governed by $\beta$-drift. This study investigates how $\beta$-drift and the resulting westward motion depend on the depth of these cyclones. Counterintuitively, shallower cyclones drift more slowly, a consequence of stronger vortex stretching. By employing a 2D quasi-geostrophic model of Jupiter’s polar regions, we constrain the cyclones’ deformation radius, a key parameter that serves as a proxy for their vertical extent, required to replicate the observed westward drift. We then explore possible vertical structures and the static stability of the poles by solving the eigenvalue problem that links the 2D model to a 3D framework, matching the constrained deformation radius. These findings provide a foundation for interpreting upcoming Juno microwave measurements of Jupiter’s north pole, offering insights into the static stability and vertical structure of the polar cyclones. Thus, by leveraging long-term motion as a novel constraint on vertical dynamics, this work sets the stage for advancing our understanding of the formation and evolution of Jupiter’s enigmatic polar cyclones.
@@
Preprint February 18, 2025
@@ 1 Introduction
Historically, our understanding of Jupiter’s atmosphere has been limited to observations of its visible cloud layer, located near the 1-bar pressure level. This limitation arises from the challenges of probing beneath the optically thick ammonia and water clouds Atreya et al. (1999); de Pater et al. (2016). Until recently, the only direct measurements of Jupiter’s subcloud layers were the localized and costly in-situ data collected by the Galileo Probe Young (1998). However, an indirect method using gravity measurements—by tracking changes in spacecraft acceleration during close flybys—has enabled insights into the large-scale flows at depth. This approach has revealed the 3D structure of zonal flows within Jupiter’s subequatorial jet streams Kaspi et al. (2018); Galanti et al. (2021); Kaspi et al. (2023) and constrained the depth of the Great Red Spot to less than $500$ km Galanti et al. (2019); Parisi et al. (2021).
Another approach to exploring Jupiter’s subcloud layers involves observations at cloud-penetrating frequencies. For example, using radio frequencies, the Earth-based Very Large Array (VLA) telescopes have probed down to pressures of approximately 8 bars de Pater et al. (2016).
@@
A more transformative development in exploring Jupiter’s subcloud atmosphere came with the Juno spacecraft. Equipped with a 6-channel Microwave Radiometer (MWR), Juno has been providing spatially resolved observations of "brightness temperatures" down to depths of approximately 240 bars Bolton et al. (2017); Janssen et al. (2017). These thermodynamic measurements are influenced both by temperature variations Fletcher et al. (2021) and the opacity of ammonia clouds Li et al. (2017). Although there is ongoing debate about the relative contributions of these factors, the observed deep patterns reveal valuable information about the vertical structure and depth of atmospheric phenomena seen at the cloud tops. For example, the MWR detected alternating zonal bands of ammonia depletion and enrichment extending to 240 bars between midlatitude jets. These patterns correspond to the upwelling and downwelling branches expected in a Ferrel-cell-like circulation Duer et al. (2021, 2023).
@@
Another significant instrument on Juno, the Jovian Infrared Auroral Mapper (JIRAM), was designed to study the aurora and atmospheric chemistry at 5-7 bar Adriani et al. (2017). During Juno’s polar orbits, as JIRAM is insensitive to half the pole being dark and can wholly observe the pole, it discovered a new phenomenon: a polar "crystal" of massive cyclones (each $\sim 5,000$ km wide) inhabiting each pole, consisting of a polar cyclone (PC) and 8 circumpolar cyclones (CPCs) at the north pole, and 5 at the south Adriani et al. (2018). The MWR instrument, designed to function optimally at the shorter distances from Jupiter’s cloud-tops during the around-equatorial skims in its original polar orbit, now enters a phase in Juno’s extended mission, where the orbit has considerably shifted, allowing MWR measurements in the north pole of Jupiter Orton et al. (2024). This enables multi-channel measurements with sufficient resolution to capture the north polar cyclones, similar to analogous previous measurements of vortices in the midlatitudes Bolton et al. (2021). This paper aims to use these novel measurements to interpret the vertical structure and extent of Jupiter’s polar cyclones.
@@
The energetic processes sustaining Jupiter’s polar cyclone configurations, while explored in prior studies O’Neill et al. (2015); Brueshaber et al. (2019); Hyder et al. (2022); Siegelman et al. (2022b, a), remain largely unconstrained. Nevertheless, substantial progress has been made in understanding the steady-state momentum balance that governs these cyclones, explaining both their stability and long-term motion. This momentum balance is closely related to the concept of $\beta$-drift—a secondary effect that causes vortices to move due to their interaction with a background vorticity gradient Sutyrin and Flierl (1994) known to influence the motion of tropical cyclones on Earth Franklin et al. (1996). The phenomenon of $\beta$-drift has been explored extensively in idealized settings, including theoretical studies Adem (1956); Smith and Ulrich (1990), numerical simulations Lam and Dritschel (2001); Gavriel and Kaspi (2023), and laboratory experiments Benzeggouta et al. (2025).
Traditionally, as suggested by the term "$\beta$-drift", studies have focused on the interaction of cyclones with $\beta$, the meridional gradient of the planetary vorticity $f$, causing a poleward-westward migration. However, in addition to $\beta$, cyclones on Jupiter’s poles exhibit mutual repulsion due to the interaction of the vorticity gradient of one cyclone with that of another. This mutual interaction allows for the stable crystal-like configuration of the cyclones Gavriel and Kaspi (2021), provided that they are surrounded by an anticyclonic annulus, commonly referred to as shielding Li et al. (2020). This repulsion also explains the oscillatory motion of the polar cyclones Gavriel and Kaspi (2022), as observed by Juno Mura et al. (2021).
@@
In this study, we investigate how the vertical structure and depth of Jupiter’s polar cyclones influence their mean westward drift. First, we employ an idealized 2D model to explore how depth—and the associated column stretching—affects $\beta$-drift on a vortex in a $\beta$-plane. Next, we perform a parameter scan using a 2D Quasi-Geostrophic (QG) polar model with Jupiter’s polar cyclones initialized in the system to constrain the QG deformation radius ($L_{d}$), a proxy for depth, such that the modeled westward drift matches observational data. Finally, we solve the eigenvalue problem posed by comparing 2D and 3D QG systems to deduce the vertical structures and static stability of the polar cyclones based on the constrained $L_{d}$.
Figure 1: Impact of Stretching on $\beta$-Drift Dynamics in a Single-Layer QG Framework. (a–b) Evolution of the zonal ($u_{\beta}$) and meridional ($v_{\beta}$) drift velocities over time for four different values of the Burger number and $b=1.2$. Dashed curves show exponential fits with amplitude $A$ and timescale $\tau$. Here, $(u_{\beta},v_{\beta})$ are non-dimensional, scaled by $\hat{\beta}V$. (c) Plots of $A$ (solid line, left axis) and $\tau$ (dashed line, right axis, in rotation periods) as functions of ${\rm Bu}$ for three values of the shape factor $b$, shown in panel (d). (d) The phase angle $\alpha$ (defined in panel e) between $u_{\beta}$ and $v_{\beta}$ vs. ${\rm Bu}$. (e) The vorticity field ($\xi_{\rm g}$, in color), streamfunction ($\psi_{\rm g}$, in contours), and velocity field ($\mathbf{u}_{\rm g}$, arrows) after five rotation periods for $\rm{Bu}=0.59$. The colorbar represents the scaled (by $\hat{\beta}V/R$) vorticity field.
2 The Impact of Vertical Depth on Vortex $\beta$-Drift
@@
A standard framework for understanding the motion of vortices due to $\beta$-drift is the barotropic quasi-geostrophic (QG) equation of a seeded vortex on a $\beta$-plane Smith and Ulrich (1990); Sutyrin and Flierl (1994); Smith (1991). In this formulation, cyclones undergo poleward-westward propagation in a planetary $\beta$, with the drift amplitude scaling with the dimensionless parameter $\hat{\beta}=\frac{\beta R^{2}}{V}$ Gavriel and Kaspi (2023), where $R$ and $V$ are the characteristic size and velocity of the cyclone. In this section, we seek to understand how the vertical extent (through stretching) modifies the classical $\beta$-drift by adding the stretching term to the QG equation.
@@
Following the derivation in Smith and Ulrich (1990) but including stretching (see supporting information for the scaling, derivation, and other details), we develop the equations in a reference frame moving with the center of the cyclone. The scaled resulting QG equation becomes (focusing on the most dominating terms)
@@
where $\xi$ and $\psi$ are the vorticity and streamfunction, respectively, and $\mathbf{u}=(u,v)$ is the velocity vector. Subscripts $\rm v$ and $\rm g$ represent, respectively, a constant background vortex flow and a "generated" field leading to the $\beta$-drift. The Burger number, ${\rm Bu}\equiv\left(\frac{L_{d}}{R}\right)^{2}$, incorporates the effect of the deformation radius, $L_{d}=\frac{\sqrt{gH}}{f_{0}}$, where $g$ is gravitational acceleration, $H$ is the layer depth, and $f_{0}$ is the Coriolis parameter. The term $-\frac{\partial}{\partial t}\left(\frac{\psi_{\rm g}}{\rm Bu}\right)$ captures the influence of stretching in the single-layer QG model.
The model is solved for vortex profiles of the form Chan and Williams (1987):
@@
where $r_{\rm c}$ is the distance from the vortex’s center, and $b$ is a shape factor for the vortex. The initial-value problem is solved using the Dedalus PDE solver Burns et al. (2020). The results from this model, as a function of ${\rm Bu}$, are shown in Fig. 1. The zonal $\beta$-drift velocity, defined as $u_{\beta}\equiv u_{\rm g}{\mid_{(x=0,y=0)}}$, displays an exponential decay, which is modeled by $A\bigl{(}1-e^{-t/\tau}\bigr{)}$ (Fig. 1a).
We further examine the dependence of the fitted maximum drift velocity ($A$) and timescale ($\tau$) on layer depth (Fig. 1c) for three values of the shape factor $b$ within a physically relevant range. The results indicate that while the drift amplitude is relatively insensitive to changes in $b$, it becomes more negative (more strongly westward) as the Burger number increases (or equivalently, as $H$ increases).
@@
However, when $H$ and ${\rm Bu}$ are finite, the formation of $\beta$-gyres (described by the term $\frac{\partial\xi_{\rm g}}{\partial t}$ in Eq. 1) is weakened, since vorticity generation now competes with stretching effects. Consequently, shallower layers experience a greater influence of stretching, which reduces the $\beta$-drift. This trend is reflected in Fig. 1c, where shallower layers correlate with stronger stretching and hence smaller $\beta$-drift velocities.
Using the Jovian values for the north and south poles of $\hat{\beta}$, $R$, and $V$ from Gavriel and Kaspi (2023), we find that matching the observed westward drift corresponds to $A=-0.46$ for the north pole and $A=-0.54$ for the south pole, marked in Fig. 1c. These values correspond to ${\rm Bu}=0.49$ for the north pole and ${\rm Bu}=0.59$ for the south, which is used in Fig. 1e to show the generated fields after five rotation periods. The influence of ${\rm Bu}$ extends beyond the reduction of the $\beta$-drift amplitude. Larger ${\rm Bu}$ values also extend the time required to reach the peak drift velocity ($\tau$ in Fig. 1c) and lead to a more zonal flow overall ($\alpha$ in Fig. 1d).
From ${\rm Bu}=(0.49,0.59)$ and $R=(811,861)$ km, the resulting deformation radii are $L_{d}=(568,658)$ km for the north and south poles, respectively. These values, however, represent simplified first estimates of $L_{d}$ that do not account for interactions among cyclones or the radial dependence of the Coriolis parameter ($f$) at Jupiter’s poles. In the next section, we consider these additional complexities to obtain a more realistic estimate of the $L_{d}$ required to explain the observed drift rates.
Figure 2: Constraining $L_{d}$ at Jupiter’s Poles Using Observed and Simulated Cyclone Drifts. (a,b) Observed trajectories of the north (a) and south (b) polar cyclones over five years Mura et al. (2022). The $\odot$ and $\otimes$ symbols mark the start and end points, respectively. The background is a JIRAM infrared image from PJ4 (Image from Adriani et al. (2018)), corresponding to the onset of these trajectories. (c,d) Simulated trajectories after five years for a model that reproduces the observed mean westward drift. The color scale and green arrows show the vorticity and velocity fields at $t=0$. (e–j) Sample model trajectories from the nested sampling routine, illustrating diverse outcomes. The numbers above the panels are the mean and standard deviation (between the cyclones) of the westward drift. The parameters used in each panel are listed below the panels. (k–n) Histogram plots of the posterior parameter distributions produced by Ultranest Buchner (2021), after sampling $10,000$ parameter sets per pole for the north (blue) and south (orange) polar cyclones. The dashed lines denote the median, and dotted lines mark the ±$\pm 1\sigma$ for the distributions. These values are written above the panels. The blue and orange points represent the values for the best model tracks (panels c,d).
3 Estimating the Deformation Radius Using a Single-Layer QG Model of the Jovian Poles
@@
with multiple cyclones prescribed by Eq. 2, with one PC and eight (five) CPCs placed initially at the observed locations during Juno’s PJ4 orbit for the north (south) poles (Fig. 2a-b). For simplicity, and due to limited observational data, we assume identical cyclones. Due to this choice and the absence of small-scale forcing, the oscillatory motion discussed by Gavriel and Kaspi (2022) is not excited. Hence, we focus on reproducing the mean westward drift.
@@
This model involves four unknown parameters: $L_{d}$, and the three cyclone profile parameters $(R,V,b)$. To explore the phase space of this model, we use the Ultranest library Buchner (2021), which employs a Bayesian nested sampling approach to identify the best-fitting solutions Buchner (2023). As a benchmark, we use the 5-year observed trajectories of the north and south polar cyclones Mura et al. (2022) (Fig. 2a-b) to calculate the average westward drift between the cyclones. This value is then used to calculate a likelihood function that evaluates the probability of a parameter set being accurate, based on the resultant model westward drift (see supporting information for details on the numerical model, likelihood function, and model sampling).
@@
We sampled $10,000$ parameter sets, integrating each simulation for seven years and calculating the resultant mean westward drift between the cyclones during the last five years (where the first 2 years act as a spin-up). Illustrative model outputs (Fig. 2 panels e–g for the north pole and panels h-j for the south) span a variety of outcomes: drifts too fast (panels e,h), too slow (panels f,i), and mergers (panels g,j). Examples of parameter sets that successfully recreate the observed drift are shown in Fig. 2c-d.
@@
Due to intrinsic parameter degeneracies, the algorithm returns a posterior distribution rather than a specific set of parameters that matches the observed westward drift. From this distribution, we obtain a more realistic estimation for the deformation radius of $L_{d}=192.41_{-38.33}^{+47.13}$ km and $L_{d}=319.15_{-33.45}^{+41.55}$ km at the $1\sigma$ confidence level for the north and south poles (Fig. 2k), respectively. The resulting posterior distributions for $R$, $V$, and $b$ (Fig. 2l–n) are consistent with plausible mean values for Jupiter’s polar cyclones and, interestingly, are very similar between the poles.

### v2

@@ Abstract
Jupiter’s poles feature striking polygons of cyclones that drift westward over time, a motion governed by $\beta$-drift (vortex motion caused by the latitudinal variation of the Coriolis force). This study investigates how $\beta$-drift and the resulting westward motion depend on the depth of these cyclones. Counterintuitively, shallower cyclones drift more slowly, a consequence of stronger vortex stretching. By employing a 2D quasi-geostrophic model of Jupiter’s polar regions, we constrain the cyclones’ deformation radius, a key parameter that serves as a proxy for their vertical extent, required to replicate the observed westward drift. We then explore possible vertical structures and the static stability of the poles by solving the eigenvalue problem that links the 2D model to a 3D framework, matching the constrained deformation radius. These findings provide a foundation for interpreting upcoming Juno microwave measurements of Jupiter’s north pole, offering insights into the static stability and vertical structure of the polar cyclones. Thus, by leveraging long-term motion as a novel constraint on vertical dynamics, this work sets the stage for advancing our understanding of the formation and evolution of Jupiter’s enigmatic polar cyclones.
@@
Preprint November 2, 2025
This article has been published in
Proceedings of the National Academy of Sciences (PNAS) (2025), 122 (44) e2503737122.
The final version is available at DOI:10.1073/pnas.2503737122.
(Received Feb 19, 2025; Revised June 24, 2025; Accepted September 16, 2025)
Key Points
  • $\beta$-drift of cyclones, linked to Jupiter’s polar cyclones’ westward drift, strengthens for deeper cyclones as vortex stretching diminishes.
  • A 2D quasi-geostrophic model estimates a deformation radius of about 220 km for the north pole and 360 km for the south pole.
  • The estimated deformation radii uncover possible vertical structures shaped by vertical stability.
Plain language summary
Jupiter’s poles feature clusters of large, long-lasting cyclones that slowly drift westward. We argue that the depth to which these cyclones extend below the cloud layers significantly influences this drift rate. By comparing computer simulations with observations, we estimate the depth of the cyclones and shed light on the stability of Jupiter’s atmosphere at high latitudes. These insights are particularly critical for Juno’s extended mission, which now allows the Microwave Radiometer (MWR) to observe these cyclones with unprecedented resolution, providing a direct test of our framework and facilitating quantitative interpretation of their vertical structure. Our results directly support this new phase of the Juno mission, offering key constraints to understand the formation, persistence, and dynamics of these giant polar storms.
@@ 1 Introduction
Figure 1: Impact of Stretching on $\beta$-Drift Dynamics in a Single-Layer QG Framework: A Single Cyclone on a $\beta$-Plane. (a–b) Evolution of the zonal ($u_{\beta}$) and meridional ($v_{\beta}$) drift velocities over time for four different values of the Burger number and $b=1.2$. Dashed curves show exponential fits with amplitude $\rm{U}_{\beta}$ and timescale $\tau$. Here, $(u_{\beta},v_{\beta})$ are non-dimensional, scaled by $\hat{\beta}V$. (c) Plots of $\rm{U}_{\beta}$ (solid line, left axis) and $\tau$ (dashed line, right axis, in rotation periods) as functions of $\rm Bu$} for three values of the shape factor $b$ (legend in panel d). The exponential parameterization becomes inadequate for ${\rm Bu}\gtrsim 5$ (see Fig. S2). (d) Solid line: Phase angle $\alpha$ (defined in panel e) between $u_{\beta}$ and $v_{\beta}$ as a function of $\rm Bu$}. Dashed line: Strength of the $\beta$-gyres, quantified as $\Delta\psi_{g}=\max(\psi_{g})-\min(\psi_{g})$, with values non-dimensionalized by $\hat{\beta}RV$. (e) The vorticity field ($\xi_{\rm g}$, in color), streamfunction ($\psi_{\rm g}$, in contours), and velocity field ($\mathbf{u}_{\rm g}$, arrows) after five rotation periods for $\rm{Bu}=0.59$ and $b=1.2$. The $\beta$-gyres appear as red and blue lobes surrounding the cyclone core. The colorbar represents the scaled (by $\hat{\beta}V/R$) vorticity field. See Movie S1 for the simulation.
@@
Historically, our understanding of Jupiter’s atmosphere has been limited to observations of its visible cloud layer, located near the 1-bar pressure level. This limitation arises from the challenges of probing beneath the optically thick ammonia and water clouds Atreya et al. (1999); de Pater et al. (2016). Until recently, the only direct measurements of Jupiter’s subcloud layers were the localized and costly in-situ data collected by the Galileo Probe Young (1998). However, an indirect method using gravity measurements (by tracking changes in the Juno spacecraft’s acceleration during close flybys) has enabled insights into the large-scale flows at depth. This approach has revealed the 3D structure of zonal flows within Jupiter’s subequatorial jet streams Kaspi et al. (2018); Galanti et al. (2021); Kaspi et al. (2023) and constrained the depth of the Great Red Spot (GRS) to less than $500$ km Galanti et al. (2019); Parisi et al. (2021).
@@
Another approach to exploring Jupiter’s subcloud layers involves observations at cloud-penetrating frequencies. For example, using radio frequencies, the Earth-based Very Large Array (VLA) telescopes have probed down to pressures of approximately 8 bars de Pater et al. (2016). The Juno Microwave Radiometer (MWR) offers an enhanced use of cloud-penetrating frequencies, leveraging its orbital proximity to achieve deeper measurements with finer spatial resolution. Equipped with six channels, the MWR has been providing spatially resolved observations of brightness temperature down to depths of approximately 240 bars Bolton et al. (2017); Janssen et al. (2017). These thermodynamic measurements are influenced both by temperature variations Fletcher et al. (2021) and the opacity of ammonia clouds Li et al. (2017). Although there is ongoing debate about the relative contributions of these factors, the observed deep patterns reveal valuable information about the vertical structure and depth of atmospheric phenomena seen at the cloud tops. For example, the MWR detected alternating zonal bands of ammonia depletion and enrichment extending to 240 bars between midlatitude jets. These patterns correspond to the upwelling and downwelling branches expected in a Ferrel-cell-like circulation Duer et al. (2021, 2023).
@@
Another key instrument on Juno, the Jovian Infrared Auroral Mapper (JIRAM), was designed to study the aurora and atmospheric chemistry at 5-7 bar Adriani et al. (2017). During Juno’s polar orbits, as JIRAM is insensitive to half the pole being dark and can wholly observe the pole, it discovered a new phenomenon: a polar "crystal" of massive cyclones (each $\sim 5,000$ km wide) inhabiting each pole, consisting of a polar cyclone (PC) and 8 circumpolar cyclones (CPCs) at the north pole, and 5 at the south Adriani et al. (2018). The MWR instrument, designed to function optimally at the shorter distances from Jupiter’s cloud-tops during the around-equatorial skims in its original polar orbit, now enters a phase in Juno’s extended mission, where the orbit has considerably shifted, allowing MWR measurements in near north pole of Jupiter Orton et al. (2024). This enables multi-channel measurements with sufficient resolution to capture the north polar cyclones, similar to analogous previous measurements of vortices in the midlatitudes Bolton et al. (2021). This study aims to provide a theoretical framework for interpreting the upcoming MWR observations of Jupiter’s north pole, offering predictions about the vertical structure and extent of the polar cyclones.
@@
The energetic processes sustaining Jupiter’s polar cyclone configurations, while explored in prior studies O’Neill et al. (2015); Brueshaber et al. (2019); Hyder et al. (2022); Siegelman et al. (2022b, a), remain largely unconstrained. Nevertheless, substantial progress has been made in understanding the principal momentum balance that governs these cyclones, explaining both their stability and long-term motion. This momentum balance is closely related to the concept of $\beta$-drift, a secondary effect that causes vortices to move due to their interaction with a background vorticity gradient Sutyrin and Flierl (1994) known to influence the motion of tropical cyclones on Earth Franklin et al. (1996). The phenomenon of $\beta$-drift has been explored extensively in idealized settings, including theoretical studies Adem (1956); Smith and Ulrich (1990), numerical simulations Lam and Dritschel (2001); Gavriel and Kaspi (2023), and laboratory experiments Firing and Beardsley (1976); Carnevale et al. (1991); Stegner and Zeitlin (1998); Flór and Eames (2002); Benzeggouta et al. (2025).
Traditionally, as suggested by the term $\beta$-drift", studies have focused on the interaction of cyclones with $\beta$, the meridional gradient of the planetary vorticity $f$, causing a poleward-westward migration through the generation of opposite vorticity anomalies (usually termed $\beta$-gyres) surrounding the core. However, in addition to $\beta$, cyclones on Jupiter’s poles exhibit mutual repulsion due to the interaction of the vorticity gradient of one cyclone with the velocity field of another. This mutual interaction allows for the stable crystal-like configuration of the cyclones Gavriel and Kaspi (2021), provided that they are surrounded by an anticyclonic annulus, commonly referred to as shielding Li et al. (2020). This repulsion also explains the oscillatory motion of the polar cyclones Gavriel and Kaspi (2022), as observed by Juno Mura et al. (2021).
@@
In this study, we investigate how the vertical structure and depth of Jupiter’s polar cyclones influence their mean westward drift. First, we employ an idealized 2D model to explore how depth, and the associated column stretching, affects $\beta$-drift on a vortex in a $\beta$-plane. Next, we perform a parameter scan using a 2D Quasi-Geostrophic (QG) polar model with Jupiter’s polar cyclones initialized in the system to constrain the QG deformation radius ($L_{d}$), a proxy for depth, such that the modeled westward drift matches observational data. Finally, we solve the eigenvalue problem posed by comparing 2D and 3D QG systems to deduce the vertical structures and static stability of the polar cyclones based on the constrained $L_{d}$.
Figure 2: Constraining $L_{d}$ at Jupiter’s Poles Using Observed and Simulated Cyclone Drifts. (a,b) Observed trajectories of the north (a) and south (b) polar cyclones over five years Mura et al. (2022). The symbols $\odot$ and $\otimes$ mark the start and end points, respectively. The background is a JIRAM infrared image from PJ4 (Image from Adriani et al. (2018)), corresponding to the onset of these trajectories. (c,d) Simulated trajectories after five years for a model that reproduces the observed mean westward drift. The color scale indicates the magnitude of the velocity vector, while the streaks represent the flow direction through line integral convolution (LIC) of the velocity field at $t=0$. (e–j) Sample model trajectories from the nested sampling routine, illustrating diverse outcomes. The numbers above the panels are the mean and standard deviation (between the cyclones) of the westward drift. The parameters used in each panel are listed below the panels. (k–n) Histogram plots of the posterior parameter distributions produced by Ultranest Buchner (2021), after sampling 15,00015{,}000 parameter sets per pole for the north (blue) and south (orange) polar cyclones. The dashed lines denote the median, and the dotted lines mark $\pm 1\sigma$ for the distributions. These values are written above the panels. Markers represent the parameter values of the successful models (panels c,d). For a corner plot of the joint posterior distributions, see Fig. S4.
The Impact of Vertical Depth on Vortex $\beta$-Drift
@@
A standard framework for understanding the motion of vortices due to $\beta$-drift is the barotropic QG equation of a seeded vortex on a $\beta$-plane Smith and Ulrich (1990); Sutyrin and Flierl (1994); Smith (1991). In this formulation, cyclones undergo poleward-westward propagation in a planetary $\beta$, with the drift amplitude scaling with the dimensionless parameter $\hat{\beta}=\frac{\beta R^{2}}{V}$ Smith (1997); Gavriel and Kaspi (2023), where $R$ and $V$ are the characteristic size and velocity of the cyclone. In this section, we seek to understand how the vertical extent (through stretching) modifies the classical $\beta$-drift by adding the stretching term to the QG equation, before treating other important complexities such as vortex-vortex interactions. Stretching, in this context, refers to the vertical elongation (or compression) of a fluid column accompanied by horizontal convergence (or divergence). In QG dynamics, such vertical motions alter the potential vorticity (PV) by concentrating or diluting the background vorticity (see Fig. S1 and the accompanying text for more intuition).
@@
Following the derivation in Smith and Ulrich (1990) but including stretching (see SI for scaling, derivation, and other details), we develop the equations in a reference frame moving with the center of the cyclone. The scaled resulting QG equation becomes (focusing on the dominant terms)
@@
where $\xi$ and $\psi$ are the vorticity and streamfunction, respectively, and $\mathbf{u}=(u,v)$ is the velocity vector. Subscripts $\rm v$ and $\rm g$ represent, respectively, a constant background vortex flow and a "generated" field leading to the $\beta$-drift. The Burger number, ${\rm Bu}\equiv\left(\frac{L_{d}}{R}\right)^{2}$, incorporates the effect of the deformation radius, $L_{d}=\frac{\sqrt{gH}}{f_{0}}$, where $g$ is gravitational acceleration, $H$ is the layer depth, and $f_{0}$ is the Coriolis parameter. The term $\frac{\partial}{\partial t}\left(\frac{\psi_{\rm g}}{\rm Bu}\right)$ captures the influence of stretching in the single-layer QG model.
The model is solved for vortex profiles of the dimensional form Chan and Williams (1987):
@@
where $r_{\rm c}$ is the distance from the vortex’s center, and $b$ is a shape factor that controls the sharpness and shielding of the vortex profile. For a more detailed analysis of the role of $b$, see Li et al. (2020). The initial-value problem is solved using the Dedalus PDE solver Burns et al. (2020). The results from this model, as a function of $\rm Bu$}, are shown in Fig. 1. The zonal $\beta$-drift velocity, defined as $u_{\beta}\equiv u_{\rm g}{\mid_{(x=0,y=0)}}$, displays an exponential relaxation with time, which is modeled by $\rm{U}_{\beta}\bigl(1-e^{-t/\tau}\bigr)$ (Fig. 1a), a parameterization that provides a good fit for $\rm{Bu}\lesssim 5$.
We further examine the dependence of the fitted maximum drift velocity ($\rm{U}_{\beta}$) and timescale ($\tau$) on layer depth (Fig. 1c) for three values of the shape factor $b$ within a physically relevant range. The results indicate that while the drift amplitude is relatively insensitive to changes in $b$, it becomes more negative (more strongly westward) as the Burger number increases (or equivalently, as $H$ increases).
@@
However, when $H$ (and thus $\rm Bu$}) is finite, the PV anomaly generated by the $\beta$-effect (the term $\hat{\beta}v_{\rm{v}}$) is divided between two parts: the formation of $\beta$-gyres (through $\frac{\partial\xi_{\rm g}}{\partial t}$) and column stretching (through $\frac{\partial}{\partial t}\frac{\psi_{\rm g}}{\rm Bu}$). As stretching has more relative significance in shallower layers, it claims a larger share of the total PV response, thereby weakening the $\beta$-gyres. This trend is reflected in Fig. 1c, where shallower layers correlate with stronger stretching and hence smaller $\beta$-drift velocities. This reduction is also evident in Fig. 1d, where the difference between the extreme streamfunction values of the $\beta$-gyres, $\Delta\psi_{g}$, becomes smaller.
Using the Jovian values for the north and south poles of $\hat{\beta}$, $R$, and $V$ from Gavriel and Kaspi (2023), we find that matching the observed westward drift corresponds to $\rm{U}_{\beta}=-0.46$ for the north pole and $\rm{U}_{\beta}=-0.54$ for the south pole, marked in Fig. 1c. These values correspond to ${\rm Bu}=0.49$ for the north pole and ${\rm Bu}=0.59$ for the south, which is used in Fig. 1e to show the generated fields after five rotation periods. The influence of $\rm Bu$} extends beyond the reduction of the $\beta$-drift amplitude. Smaller $\rm Bu$} values also reduce the time required to reach the peak drift velocity ($\tau$ in Fig. 1c) and lead to a more zonally directed drift overall ($\alpha$ in Fig. 1d).
From ${\rm Bu}=(0.49,0.59)$ and $R=(811,861)$ km Gavriel and Kaspi (2023), the resulting deformation radii are $L_{d}=(568,658)$ km for the north and south poles, respectively. These values should not be interpreted as definitive physical estimates, but rather as outcomes of an idealized model that omits vortex–vortex interactions and the polar variation of $f$. Instead, this simplified framework is meant to provide insight into the role of stretching alone in modulating $\beta$-drift. In the next section, we consider these additional complexities to derive more realistic values of $L_{d}$ needed to match the observed drift rates.
Estimating the Deformation Radius Using a Single-Layer QG Model of the Jovian Poles
@@
with multiple cyclones prescribed by Eq. 2, with one PC and eight (five) CPCs placed initially at the observed locations during Juno’s PJ4 orbit for the north (south) poles (Fig. 2a-b). For simplicity and because of limited observational data, we assume identical cyclones. Due to this choice and the absence of small-scale forcing, the oscillatory motion discussed by Gavriel and Kaspi (2022) is not excited. Hence, we focus on reproducing the mean westward drift. While zonal jets could, in principle, influence the drift, no such flows have been observed in the vicinity of the cyclones Rogers et al. (2022). Moreover, JunoCam imagery is consistent with a two-dimensional, isotropic turbulent regime Orton et al. (2017), and thus zonal flows are not included in this model.
@@
This model involves four unknown parameters: $L_{d}$, and the three cyclone profile parameters $(R,V,b)$. To explore the phase space of this model, we use the Ultranest library Buchner (2021), which employs a Bayesian nested sampling approach to identify the best-fitting solutions Buchner (2023). As a benchmark, we use the 5-year observed trajectories of the north and south polar cyclones Mura et al. (2022) (Fig. 2a-b) to calculate the average westward drift between the cyclones. This value is then used to calculate a likelihood function that evaluates the probability of a parameter set being accurate, based on the resultant model westward drift (see SI for details on the numerical model, likelihood function, and model sampling).
@@
We sampled 15,00015{,}000 parameter sets for each pole, integrating each simulation for seven years and calculating the resultant mean westward drift between the cyclones during the last five years (where the first 2 years act as a spin-up). Illustrative model outputs (Fig. 2 panels e–g for the north pole and panels h-j for the south) span a variety of outcomes: drifts too fast (panels e,h), too slow (panels f and i), eastward (panel g), and mergers (panel j). Examples of parameter sets that successfully recreate the observed drift are shown in Fig. 2c-d.
@@
Due to intrinsic parameter degeneracies, the algorithm returns a posterior distribution rather than a specific set of parameters that matches the observed westward drift. From this distribution, we obtain a more realistic estimate for the deformation radius of $L_{d}=225.58_{-46.93}^{+83.94}$ km and $L_{d}=363.69_{-44.33}^{+74.50}$ km at the $1\sigma$ confidence level for the north and south poles (Fig. 2k), respectively. The resulting posterior distributions for $R$, $V$, and $b$ (Fig. 2l–n) align well with plausible mean values for Jupiter’s polar cyclones, as estimated from cloud feature tracking at the poles Grassi et al. (2018).